# Language Model Planning from an Information Theoretic Perspective

### v1

@@ 4 Conclusion
Drawing on a VQ-VAE, we develop an information‑theoretic pipeline for compressing hidden‑state trajectories into discrete codes and use them to compute MI between different parts of a LM’s computation. This coarse‑grained view allows us to examine how much the prefix internal computations inform future prediction states, whether they preserve alternative continuations, and which blocks of history a decoder‑only transformer relies on. Empirically, we compare training with standard NTP and MTP losses across a synthetic CFG task, PF tasks, and natural text. Our findings show that planning behavior varies with the task. In the CFG task, nMI between the prefix and decision state quickly decays, suggesting myopic computations. In path‑finding, however, the prefix retains strong information about future steps and alternative correct paths, indicating branch awareness. We also observe a recency effect: decision states depend mainly on the final layers and most recent block, though earlier activations still carry some influence, with most of the dependence coming from the last prefix token. Taken together, these results suggest that LMs perform planning, but the extent and nature of that planning are contingent on the task and training objective. Our VQ-VAE-based MI framework provides an automated way to investigate such behaviors. Future work could extend this analysis to chain‑of‑thought‑style training and RL‑fine‑tuned LMs, and investigate how architectural modifications can encourage planning.
@@ Acknowledgements.
This work was partially supported by the US National Science Foundation under grants CCF 2045694, CNS-2112471, CPS-2111751, and ONR N00014-23-1-2149 to GJ. In addition, the work was supported by NSF Grants 2154171, CAREER Award 2339112, CMU CyLab Seed Funding. This work used PSC Bridges-2 GPU at Pittsburgh Supercomputing Center through allocation CIS250226 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by US National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
@@ A.2 Training and losses of VQ-VAE
We train each VQ-VAE on hidden state blocks $G_{S}$ extracted from a fixed, pretrained LM $\mathcal{M}$ over datasets used in our analyses. For each mini-batch, we sample sequences ${\mathbf{x}}_{1:T}$, feed $\mathcal{M}$ with inputs to get hidden activations $h^{\ell}_{t}$, assemble the target blocks $G_{S}$ according to the experiment. Then, we optimize encoder $E$, codebook $\{e_{k}\}_{k=1}^{K}$, and decoder $D$ while keeping $M$ frozen. This follows the standard VQ-VAE pipeline [57] adapted to variable-shape transformer hidden states.
@@
and reconstruct $\widehat{G}_{S}=D(\tilde{r}_{S})$ with straight-through gradients for the quantizer [57]. The loss combines four parts
@@
Quantization and commitment follow [57] with the straight-through estimator
@@
which spreads codebook vectors and yields more discriminative high-level codes, consistent with our use of codes as discussed in main text § 2. To avoid mode collapse of VQ-VAE [66, 68], we further uses an anti-collapse entropy term which employs soft assignments computed from distances to the full codebook. Let
@@
VQ-VAE training can suffer from codebook collapse [66, 68]. We adopt three complementary heuristics. (i) Dead-code reset: codes that have not been the argmin $k^{\star}$ for a fixed window are reinitialized to random encoder outputs from the current batch. (ii) Codebook dropout: with probability $0.1$ we temporarily mask a random subset of codebook vectors when forming nearest-neighbor distances in the forward pass of training, which encourages redundancy avoidance and better exploration. (iii) Entropy regularization: the $\mathcal{L}_{\text{ent}}$ term above explicitly pushes assignments away from degenerate peaks.
@@
A.3 A Validation on a representative experiment
To validate our use of VQ-VAE-derived codes for estimating between computations, we define discrete variables $(A,B)$. Let $A$ be uniform on a domain $\mathcal{A}$ with $|\mathcal{A}|=1024$, and let
@@ B.1 Loss functions
To align internal states with short-horizon futures, we equip the model with $\Gamma$ output heads [18]. Head $\gamma\in\{1,\dots,\Gamma\}$ is implemented as a transformer layer $f_{h}^{(\gamma)}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ that takes $h_{\scriptscriptstyle t}^{\scriptscriptstyle L}$} as input and is followed by the same shared unembedding $U$:
@@ B.2.1 Context-free grammar data generation details
The specific formal details of our CFG are not central to the claims of this paper. We use it only as a controlled source of long, purely syntactic sequences. Readers who want a formal presentation and broader context can consult the CFG-based data generation reference [4], on which our CFG data is based on. Here we provide the details of our specific CFG parallel to the naming convention in [4].
@@ B.2.4 How the generated data are used to train the language models
B.5 Details of the history in the plan experiment
@@
We provide the results in Figures 9&10.
Figure 9: Heatmap of normalized mutual information over block and layer indices for $\tau=0,1,2$. Please refer to the caption of 3.
Figure 10: Heatmap of normalized mutual information over block and layer indices for $\tau=3,4,5$. Please refer to the caption of 3.
@@ Appendix C Related work in detail
In order to improve LMs’ reasoning and planning abilities, researchers have developed scaffolding and augmentation techniques, including Chain-of-Thought prompting [64], planning tokens [63, 46], and structured inference methods such as Tree-of-Thoughts and Graph-of-Thoughts [65, 7]. In parallel, hybrid systems that integrate LMs with symbolic planners or external tools have achieved state-of-the-art performance in embodied and tool-use domains [67, 60, 48]. Despite these advances, recent studies highlight that significant challenges remain, and current approaches to LM planning still fall short of fully addressing complex reasoning and decision-making tasks. For example, [32] show that models can produce conflicting answers under logically related prompts despite local plausibility, which shows lack of long horizon planning, although short-horizon planning works fine; [2] reveal discrepancies when asking a model what is correct vs what is incorrect; and [45] demonstrate that even repeated queries yield inconsistent outputs. Beyond inconsistency, some studies find that out-of-the-box models struggle with even simple planning tasks [37]. Complementing these empirical findings, theoretical analyses suggest that autoregressive models may face fundamental limits in their planning ability [62]. Taken together, these findings underscore that understanding whether and how planning arises in LMs is not only an open empirical challenge, but also central to both their interpretability and the principled design of future model architectures.
@@
Some interpretability (explainability) methods consider the LM as black box, and design tasks or benchmarks to gauge reasoning, robustness, or generalization abilities [51, 30]. In contrast to them, mechanistic interpretability seeks to reverse-engineer transformer computations into human-understandable parts, treating the residual stream as the main information pathway and attention heads as separate components that pass information along [15]. Using this approach, researchers have explained key phenomena in LMs: in-context learning can arise from a characteristic two-head circuit that appears during a training “phase change,” with causal support from perturbation studies [40]; many neurons are polysemantic, but sparse autoencoders (SAEs) can replace them with more interpretable, monosemantic features that enable finer-grained causal analysis [12, 8]; models implicitly encode structural linguistic properties such as syntax [25]. Building on earlier approaches, transcoders, approximate dense MLPs with wider sparse layers, separating circuits into input-invariant and input-dependent components, and matching or surpassing SAEs in sparsity, faithfulness, and interpretability [14]. Frontier-scale case studies use cross-layer transcoders to trace multi-step reasoning and other behaviors, illustrating how these tools can audit mechanisms, not just features, in modern LLMs [33]. Overall, mechanistic interpretability offers concrete tools for uncovering how LLMs compute, though these methods are still developing.
@@
Beyond empirical tools, mathematically grounded perspectives also illuminate transformer/LLM interpretability: automata-theoretic analyses show self-attention can implement finite-state algorithms [34]; optimization views prove in-context learning corresponds to (preconditioned) gradient steps [3] and transformers are mutual interaction learners [56]; mean-field theory gives global convergence guarantees at scale [16].
@@
A different line of interpretability work views LLM hidden states as structured representations that can be "probed", probing refers to training lightweight models, often linear classifiers, to read out specific information from hidden states in order to test what the model represents internally. Using probing, researchers have shown that transformer hidden states encode structured belief-state and world-model–like information [47, 21, 24]. Other works demonstrate that a single hidden state can carry information about multiple future tokens [41] and that probing can reveal the underlying algorithms LLMs use to solve tasks [4].

### v2

@@ 4 Conclusion
Drawing on a VQ-VAE, we develop an information-theoretic pipeline that compresses hidden-state trajectories into discrete codes and uses them to compute MI across an LM’s computation. This lens measures (i) how prefix computations inform future decision states, (ii) whether models retain information about alternative continuations, and (iii) how decision-relevant information is distributed across layers and prefix blocks. We evaluate models trained with standard NTP versus MTP losses on a synthetic CFG task, PF tasks, and natural text; we find no consistent differences between NTP- and MTP-trained models. The results are strongly task-dependent: in CFG, nMI between the prefix and the decision state decays quickly, consistent with largely myopic computation; in path-finding, the prefix retains substantial information about later steps and alternative correct paths, indicating stronger branch awareness. Across settings, we observe a recency effect, decision states depend most on late layers and recent blocks, while earlier activations remain measurably informative. Overall, LMs exhibit internal planning, but its extent and form vary with the task and training objective; our VQ-VAE-MI framework provides an automated way to study these behaviors. Future work could extend this analysis to reasoning models and test architectural changes that promote planning.
@@ Acknowledgements.
This work was partially supported by the US National Science Foundation under grants CCF 2045694, CNS-2112471, CPS-2111751, ONR N00014-23-1-2149 to GJ and US Department of Energy under grant DESC0025652 to CJW. In addition, the work was supported by Pennsylvania Infrastructure Technology Alliance, NSF Grants 2154171, CAREER Award 2339112, NSF Award 2512805, CMU CyLab Seed Funding to GQ. This work used PSC Bridges-2 GPU at Pittsburgh Supercomputing Center through allocation CIS250226 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by US National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
Reproducibility statement.
We describe our method in § 2, with further details provided in the relevant appendices. Our experimental analyses are presented in § 3, with additional details likewise included in the appendices. To ensure reproducibility, we provide the full code, tools to acquire the datasets, and instructions in the supplementary material. In addition, we include a README file that explains how to run each experiment separately.
@@ A.2 Training and losses of VQ-VAE
We train each VQ-VAE on hidden state blocks $G_{S}$ extracted from a fixed, pretrained LM $\mathcal{M}$ over datasets used in our analyses. For each mini-batch, we sample sequences ${\mathbf{x}}_{1:T}$, feed $\mathcal{M}$ with inputs to get hidden activations $h^{\ell}_{t}$, assemble the target blocks $G_{S}$ according to the experiment. Then, we optimize encoder $E$, codebook $\{e_{k}\}_{k=1}^{K}$, and decoder $D$ while keeping $M$ frozen. This follows the standard VQ-VAE pipeline (Van Den Oord et al., 2017) adapted to variable-shape transformer hidden states.
@@
and reconstruct $\widehat{G}_{S}=D(\tilde{r}_{S})$ with straight-through gradients for the quantizer (Van Den Oord et al., 2017). The loss combines four parts
@@
Quantization and commitment follow Van Den Oord et al. (2017) with the straight-through estimator
@@
which spreads codebook vectors and yields more discriminative high-level codes, consistent with our use of codes as discussed in main text § 2. To avoid mode collapse of VQ-VAE (Zhao et al., 2024; Zhu et al., 2025), we further uses an anti-collapse entropy term which employs soft assignments computed from distances to the full codebook. Let
@@
VQ-VAE training can suffer from codebook collapse (Zhao et al., 2024; Zhu et al., 2025). We adopt three complementary heuristics. (i) Dead-code reset: codes that have not been the argmin $k^{\star}$ for a fixed window are reinitialized to random encoder outputs from the current batch. (ii) Codebook dropout: with probability $0.1$ we temporarily mask a random subset of codebook vectors when forming nearest-neighbor distances in the forward pass of training, which encourages redundancy avoidance and better exploration. (iii) Entropy regularization: the $\mathcal{L}_{\text{ent}}$ term above explicitly pushes assignments away from degenerate peaks.
@@
A.3 VQ-VAE Training Setup Across Experiments
Across all experiments, each VQ-VAE is trained once per dataset/task and representation type. This design ensures that the quantization mechanism remains fixed within each study, avoiding confounding effects that would arise from retraining VQ-VAE models at different mutual information estimates. In the planning horizon experiments, each underlying dataset is equipped with its own hierarchical pair of VQ-VAEs ($\texttt{VQ-VAE}_{1}$ and $\texttt{VQ-VAE}_{2}$) for quantizing prefix hidden-state blocks, along with its own last-layer VQ-VAE. Thus, while the architectural pattern is shared, the models themselves are trained separately for each task. In the branching experiments, we again train $\texttt{VQ-VAE}_{1}$ and $\texttt{VQ-VAE}_{2}$ for prefix hidden states, together with a dedicated path-sequence VQ-VAE. In the history experiments, we train one VQ-VAE for prefix blocks and one for last-layer OpenWebText representations. Once trained, all VQ-VAE models are reused for every mutual-information estimation within their respective experimental setting.
With respect to training budgets, the VQ-VAE models require only moderate optimization steps. In the planning horizon experiments, VQ-VAE,1 and VQ-VAE,2 (used for prefix hidden states) are trained for 15,000 optimization steps each, corresponding to 0.96M sequences in total, while the last-layer VQ-VAE is trained for 10,000 steps (0.32M sequences). The path-finding version of this experiment uses the same budgets. These settings apply uniformly to both next-token and multi-token prediction evaluations. In the branching experiments, VQ-VAE,1 and VQ-VAE,2 (for prefix states) are trained for 20,000 (0.64M sequences) and 15,000 steps (0.48M sequences), respectively, and the path-sequence VQ-VAE is trained for 25,000 steps (0.80M sequences). Overall, this experiment uses between 0.32M and 0.96M sequences per VQ-VAE depending on the role. In the history experiments, we train two VQ-VAEs, one for prefix hidden-state blocks and one for last-layer hidden states, each for 10,000 optimization steps, corresponding to 0.64M sequences per model.
Choosing hyperparameters for our VQ-VAE models requires some initial effort, but not because the models are inherently sensitive. Instead, the difficulty arises from the fact that our VQ-VAE objective includes several additional components beyond the conventional formulation, such as cosine-push regularization, entropy-based loss, and codebook-reset mechanisms. These additions interact in nontrivial ways, and as a result the first working configuration for the first experiment takes time to establish simply because there is no pre-existing recipe for our variant of the objective. Once this initial configuration is found, however, the process becomes easier. We consistently find that once a reasonable setting is chosen for one model, closely related values work across all others with little to no retuning. The only exception is the codebook size. Starting from very small values, increasing the codebook size improves VQ-VAE performance, but only up to two possible limits. The first limit is model size. If the VQ-VAE is too small, increasing the codebook size provides little benefit, and making the model larger restores the usefulness of larger codebooks. The second limit is the inherent complexity of the task. After a certain point, increasing the codebook size and the model size no longer leads to meaningful improvements. For these reasons, our experiments use codebook sizes ranging from 64 (for quantizing a single token in a single-layer hidden state) up to 1024 (for quantizing all prefix hidden states). For completeness and reproducibility, all hyperparameters can be accessed in the configurator files provided with the code.
A.4 Validation on representative experiments
To validate our use of VQ-VAE-derived codes for estimating between computations, we conduct three experiments below with different difficulty levels.
A.4.1 Validation experiment with known mutual information
We define discrete variables $(A,B)$. Let $A$ be uniform on a domain $\mathcal{A}$ with $|\mathcal{A}|=1024$, and let
@@ A.4.1 Validation experiment with known mutual information
A.4.2 Validation experiment (harder setting with multiple similar prefixes for each label)
We construct a more challenging variant of App. A.4.1. For each $a\in\mathcal{A}$ we now create a set of $16$ surrogate hidden-state tensors in $\mathbb{R}^{10\times 11\times 768}$. Concretely, we take a random length-$10$ prefix and generate $15$ additional prefixes by changing a single token at a random index, then run the trained LM to collect the corresponding hidden state block outputs. Similarly, for each $b$ we produce $16$ final-layer hidden states in $\mathbb{R}^{768}$ using the same single-token perturbation strategy. Therefore, this experiment consists of $2^{14}$ possible hidden state block $G_{A}$ and $2^{14}$ possible final-layer hidden state $G_{B}$. This setting reflects families of inputs that differ at one token yet retain essentially the same semantics, which is exactly the kind of fine detail our coarse summaries should compress away. To further increase difficulty, we add Gaussian noise during VQ-VAE training on hidden states. We then compute normalized mutual information between the learned coarse variables, treating the $16$ surrogates for a fixed label as equiprobable when forming empirical distributions. The results in Figure 6 show that for every codebook size $K$, the codes preserve the ground-truth ordering across $p_{v}$, and $\mathcal{I}\left(Z_{A};Z_{B}\right)$ tightens toward the reference as $K$ increases.
Figure 6: Harder validation experiment with similar prefixes for each label. Each label maps to $16$ surrogates created by one-token edits. Normalized $\mathcal{I}\left(Z_{A};Z_{B}\right)$ remains order-consistent with the reference-process MI across $K$, despite injected noise and within-class variability.
A.4.3 Validation experiment (hardest setting with multiple unrelated prefixes)
We now remove within-class similarity entirely. For each $a\in\mathcal{A}$ we sample $16$ independent length-$10$ token sequences uniformly at random and collect their $\mathbb{R}^{10\times 11\times 768}$ hidden-state tensors. For each $b$ we likewise sample $16$ independent sequences and extract the corresponding $\mathbb{R}^{768}$ final-layer states. Thus, the $16$ surrogates that share a label need not be close in vector space. This setting emulates many-to-one mappings from highly diverse inputs to the same high-level code. This setting resembles natural language settings with dissimilar texts having similar meanings. Again, this experiment consists of $2^{14}$ possible hidden state block $G_{A}$ and $2^{14}$ possible final-layer hidden state $G_{B}$.
We repeat the same pipeline and probability treatment as in App. A.4.2, estimating normalized MI between the learned coarse variables while assuming the $16$ surrogates per label are equiprobable. Figure 7 reports the results. Although this regime poses $16$ equally likely outcomes per input and eliminates structure exploitable by local similarity, our method still recovers normalized MI that preserves the correct ordering across $p_{v}$, with improved agreement at larger $K$. This regime would be quite challenging for a probing-type method to recover any meaningful information due to one-to-many mapping with totally unrelated targets.
Figure 7: Hardest validation experiment with fully independent surrogates. Despite the lack of within-class proximity, normalized $\mathcal{I}\left(Z_{A};Z_{B}\right)$ preserves the ordering induced by the reference-process MI, and improves as the codebook grows.
@@ B.1 Loss functions
To align internal states with short-horizon futures, we equip the model with $\Gamma$ output heads Gloeckle et al. (2024). Head $\gamma\in\{1,\dots,\Gamma\}$ is implemented as a transformer layer $f_{h}^{(\gamma)}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ that takes $h_{\scriptscriptstyle t}^{\scriptscriptstyle L}$} as input and is followed by the same shared unembedding $U$:
@@ B.2.1 Context-free grammar data generation details
The specific formal details of our CFG are not central to the claims of this paper. We use it only as a controlled source of long, purely syntactic sequences. Readers who want a formal presentation and broader context can consult the CFG-based data generation reference Allen-Zhu (2024), on which our CFG data is based on. Here we provide the details of our specific CFG parallel to the naming convention in Allen-Zhu (2024).
@@ B.2.4 How the generated data are used to train the language models
B.5 Details of the information in the computational history experiment
In this section we provide details of the information in the computational history experiment.
@@ B.5 Details of the information in the computational history experiment
We provide the results in Figures 12&13.
Figure 12: Heatmap of normalized mutual information over block and layer indices for $\tau=0,1,2$. Please refer to the caption of 4.
Figure 13: Heatmap of normalized mutual information over block and layer indices for $\tau=3,4,5$. Please refer to the caption of 4.
@@ Appendix C Related work in detail
In order to improve LMs’ reasoning and planning abilities, researchers have developed scaffolding and augmentation techniques, including Chain-of-Thought prompting (Wei et al., 2022), planning tokens (Wang et al., 2024c; Sel et al., 2025), and structured inference methods such as Tree-of-Thoughts and Graph-of-Thoughts (Yao et al., 2023; Besta et al., 2024). In parallel, hybrid systems that integrate LMs with symbolic planners or external tools have achieved state-of-the-art performance in embodied and tool-use domains (Zhao et al., 2023; Wang et al., 2024a; Shen et al., 2023). Despite these advances, recent studies highlight that significant challenges remain, and current approaches to LM planning still fall short of fully addressing complex reasoning and decision-making tasks. For example, Lin et al. (2025) show that models can produce conflicting answers under logically related prompts despite local plausibility, which shows lack of long horizon planning, although short-horizon planning works fine; Ahn and Yin (2025) reveal discrepancies when asking a model what is correct vs what is incorrect; and Saxena et al. (2024) demonstrate that even repeated queries yield inconsistent outputs. Beyond inconsistency, some studies find that out-of-the-box models struggle with even simple planning tasks (Momennejad et al., 2023). Complementing these empirical findings, theoretical analyses suggest that autoregressive models may face fundamental limits in their planning ability (Wang et al., 2024b). Taken together, these findings underscore that understanding whether and how planning arises in LMs is not only an open empirical challenge, but also central to both their interpretability and the principled design of future model architectures.
@@
Some interpretability (explainability) methods consider the LM as black box, and design tasks or benchmarks to gauge reasoning, robustness, or generalization abilities (Srivastava et al., 2023; Liang et al., 2022). In contrast to them, mechanistic interpretability seeks to reverse-engineer transformer computations into human-understandable parts, treating the residual stream as the main information pathway and attention heads as separate components that pass information along (Elhage et al., 2021). Using this approach, researchers have explained key phenomena in LMs: in-context learning can arise from a characteristic two-head circuit that appears during a training “phase change,” with causal support from perturbation studies (Olsson, 2022); many neurons are polysemantic, but sparse autoencoders (SAEs) can replace them with more interpretable, monosemantic features that enable finer-grained causal analysis (Cunningham et al., 2023; Bricken et al., 2023); models implicitly encode structural linguistic properties such as syntax (Hewitt and Liang, 2019). Building on earlier approaches, transcoders, approximate dense MLPs with wider sparse layers, separating circuits into input-invariant and input-dependent components, and matching or surpassing SAEs in sparsity, faithfulness, and interpretability (Dunefsky et al., 2024). Frontier-scale case studies use cross-layer transcoders to trace multi-step reasoning and other behaviors, illustrating how these tools can audit mechanisms, not just features, in modern LLMs (Lindsey et al., 2025). Overall, mechanistic interpretability offers concrete tools for uncovering how LLMs compute, though these methods are still developing.
@@
Beyond empirical tools, mathematically grounded perspectives also illuminate transformer/LLM interpretability: automata-theoretic analyses show self-attention can implement finite-state algorithms (Liu et al., 2023); optimization views prove in-context learning corresponds to (preconditioned) gradient steps (Ahn et al., 2023) and transformers are mutual interaction learners (Ustaomeroglu and Qu, 2025); mean-field theory gives global convergence guarantees at scale (Gao et al., 2024).
@@
A different line of interpretability work views LLM hidden states as structured representations that can be ”probed”, probing refers to training lightweight models, often linear classifiers, to read out specific information from hidden states in order to test what the model represents internally. Using probing, researchers have shown that transformer hidden states encode structured belief-state and world-model–like information (Shai et al., 2024; Gurnee and Tegmark, 2024; Hazineh et al., 2023). Other works demonstrate that a single hidden state can carry information about multiple future tokens (Pal et al., 2023) and that probing can reveal the underlying algorithms LLMs use to solve tasks (Allen-Zhu, 2024).